# Constructing new probability distributions on the unit interval

### v1

@@ Abstract
In this paper, we introduce a novel method for constructing probability distributions on the unit interval by exploiting the non-injective transformation defined by the ratio of two positive random variables, $X$ and $Y$. For simplicity and tractability, we focus on independent random variables $X$ and $Y$ following gamma distributions. We derive some results including laws, density functions, quantile function and closed-form expressions for the moments, and discuss maximum likelihood estimation.
@@ 1 Introduction
Probability models with suppport on the unit interval are usefull in many fields, including econometrics (Cribari-Neto and Zeileis,, 2010), natural sciences Geissinger et al., (2022), and reliability Cross et al., (2006), among others. Traditionally, the beta (Ferrari and Cribari‐Neto,, 2004) and Kumaraswamy (Kumaraswamy,, 1980) distributions have been the workhorses for modeling such data, due to their flexibility and tractability. More recently, alternative unit‐interval families have emerged, see, for example, Vila et al., 2024c ; Vila et al., 2024b , to address features like bimodality.
The beta distribution is constructed by the ratio of two positive random variables, that is, when $X,Y\sim\mathrm{Gamma}(\alpha_{i},\lambda)$ are independent and identically scaled, then
@@
yielding the familiar beta model. In this work, we introduce two new families of unit‐interval distributions, defined by non‐injective transformations of the ratio $X/(X+Y)$. Writing
[…]
The remainder of this paper is organized as follows. In Section 2, we introduce the new unit‐interval families via the non‐injective transformations $S_{r}$ and $T_{r}$ of the gamma ratio. In Section 3, we derive their distributional laws in terms of the underlying $X/(X+Y)$ density. In Section 4, we obtain closed‐form expressions for the density and cumulative distribution functions of $W$ and $Z$. In Section 5, we present the quantile function. In Section 6, we develop closed‐form formulae for selected raw moments under special parameter regimes. In Section 7, we formulate maximum likelihood estimation, derive the likelihood equations, and discuss numerical solution strategies. Finally, in Section 8, we provide some concluding remarks.
@@ 2 The unit probabilistic models
[…]
${X\over X+Y}\sim\text{Beta}(\alpha_{1},\alpha_{2}).$
[…]
@@
Note that both transformations above have a zero at the point $x=1/2$ and assume the value of 111 at the points $x=0$ and $x=1$, see Figure 1.
Figure 1: Plots of the transformations $S_{r}$ and $T_{r}$ for different choices of the parameter $r$.
@@
$W\equiv S_{r}\left({X\over X+Y}\right)=\left({Y-X\over X+Y}\right)^{2r},$ (3)
$Z\equiv T_{r}\left({X\over X+Y}\right)=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}},$ (4)
@@
Remark 2.2.
The value of $r=1/2$ in formulas (3) and (4) is special, because for this value the variables $W$ and $Z$ take the forms
$W={|X-Y|\over X+Y}\quad\text{and}\quad Z=1-{\sqrt{XY}\over{1\over 2}(X+Y)},$ (5)
@@ 3 Deriving the laws of $W$ and $Z$
$\mathbb{P}\left(W\in B\right)=\mathbb{P}\left(S_{r}\left({X\over X+Y}\right)\in B\right)$ =($=\mathbb{P}\left(\left\{\omega\in\Omega:{X(\omega)\over X(\omega)+Y(\omega)}\in S_{r}^{-1}(B)\right\}\right)$
$=\int_{S_{r}^{-1}(B)}f_{X\over X+Y}(u){\rm d}u,$ (6)
@@ 4 Distribution and density functions of $W$ and $Z$
Figures 2-7 display a range of shapes for the densities of $W$ and $Z$.
Figure 2: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=0.3$.
@@
Figure 4: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=0.5$.
@@
Figure 7: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=3$.
@@
It is worth noting that in the case $\lambda_{1}=\lambda_{2}\equiv\lambda$, the densities $f_{W}$ and $f_{Z}$ presented in Remark 4.1 do not depend on $\lambda$, consistent with the fact that $W$ and $Z$, given in (3) and (4), respectively, are derived from scale-invariant transformations. That is, for $X\sim\text{Gamma}(\alpha_{1},\lambda)$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda)$ we have $\lambda X\sim\text{Gamma}(\alpha_{1},1)$ and $\lambda Y\sim\text{Gamma}(\alpha_{2},1)$, and consequently, the random variables $W$ and $Z$ satisfy the identities
$W=\left({Y-X\over X+Y}\right)^{2r}=\left({\lambda Y-\lambda X\over\lambda X+\lambda Y}\right)^{2r},$
$Z=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}}=1-{(\lambda X\lambda Y)^{r}\over\displaystyle\left[{1\over 2}(\lambda X+\lambda Y)\right]^{2r}},$
@@ 6 Closed-form expressions for the $n$th raw moments of $W$ and $Z$
$\mathbb{E}(W^{n})$ $={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}w^{n+{1\over 2r}-1}G(w_{r}){\rm d}w$
and
$\mathbb{E}(Z^{n})$ $={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}{z^{n}(1-z)^{{1\over r}-1}G(z_{r})\over\sqrt{1-(1-z)^{1\over r}}}{\rm d}z,$
where $w_{r}$ and $z_{r}$ are as given in (11) and (14), respectively, and $G(\cdot)$ is as in (18). Note that the above integrals are well-defined due to the support of $W$ and $Z$ in (0,1)(0,1)( 0 , 1 ), but numerical methods are necessary for their evaluation due to analytical complexity.
@@
6.1 The 111st raw moment of $W$: case $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$
Under the conditions $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$, in this part, we obtain (more precisely, in (25) and Remark 6.2) straightforward closed-form expressions for the 111st raw moment (expected value) of
@@ 6.1 The 111st raw moment of $W$: case $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$
$\mathbb{E}(W)=\mathbb{E}\left[\frac{|X-Y|}{X+Y}\right]$ $=\mathbb{E}\left[|X-Y|\int_{0}^{\infty}\exp\left\{-\left(X+Y\right)x\right\}{\rm d}x\right]$
$=\int_{0}^{\infty}\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]{\rm d}x,$ (20)
@@
$\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]$ $=\int_{0}^{\infty}\left[\int_{u}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$=\int_{0}^{\infty}\left[\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
@@
$\int_{0}^{\infty}\Bigg{[}$ $\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\Bigg{]}\exp(-ux)=\int_{0}^{\infty}\exp(-ux){\rm d}F_{X}(u)\int_{0}^{\infty}v\exp(-vx){\rm d}F_{Y}(v)$
$-\int_{0}^{\infty}u\exp(-ux){\rm d}F_{X}(u)\int_{0}^{\infty}\exp(-vx){\rm d}F_{Y}(v)$
$=\mathbb{E}\left[\exp(-Xx)\right]\mathbb{E}\left[Y\exp(-Yx)\right]-\mathbb{E}\left[X\exp(-Xx)\right]\mathbb{E}\left[\exp(-Yx)\right]$
@@
$+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$=2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)$
$-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$
@@
$+2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)$
$-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$ (22)
@@
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}.$ (23)
@@
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x.$ (24)
@@
$\mathbb{E}(W)$ $={\alpha_{2}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2}+1,\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)$
$-{\alpha_{1}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2},\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)$
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x$

### v2

# On classes of distributions on the unit interval: structural properties and application to inequality data

@@ Abstract
Probability distributions defined on the unit interval are widely used in fields ranging from econometrics to reliability studies. Traditional models such as the beta and Kumaraswamy distributions are well-established due to their flexibility and tractability. In this paper, we introduce two novel families of unit-interval distributions derived via non-injective transformations of the gamma ratio. These transformations, denoted $S_{r}$ and $T_{r}$, allow the construction of new random variables with support on $(0,1)$ and admit simple closed-form expressions for their densities when the underlying variables are independent gamma distributed. Notably, for $r=1/2$, these constructions yield sample-based estimators of the Gini and Atkinson indices, establishing a direct link with classical inequality measures. We derive the distributional laws, cumulative distribution functions, quantile functions, and raw moments, and discuss maximum likelihood estimation for the proposed models. A Monte Carlo simulation study is conducted to assess the finite sample behavior of the maximum likelihood estimators under different parameter configurations. An application to cross-country Gini index data illustrates the flexibility and practical relevance of the proposed distributions in modeling real inequality indicators.
@@ 1 Introduction
Probability models with support on the unit interval are useful in many fields, including econometrics (Cribari-Neto and Zeileis,, 2010), natural sciences Geissinger et al., (2022), and reliability (Cross et al.,, 2006), among others. In the context of economic analysis, several key indicators are naturally constrained to $(0,1)$, most notably measures of income inequality such as the Gini index (Gini,, 1936). This coefficient is routinely used to summarize the degree of income concentration across countries.
Traditionally, the beta (Ferrari and Cribari‐Neto,, 2004) and Kumaraswamy (Kumaraswamy,, 1980) distributions have been the workhorses for modeling bounded continuous data, due to their flexibility and tractability. More recently, alternative unit–interval families have emerged, see, for example, Vila et al., 2024c ; Vila et al., 2024b , to address features such as bimodality and complex shapes. Despite these advances, most existing models are essentially phenomenological, offering limited theoretical connection with the stochastic mechanisms underlying classical inequality measures.
The beta distribution itself admits an interpretation as the ratio of two positive random variables. Indeed, when $X,Y\sim\mathrm{Gamma}(\alpha_{i},\lambda)$ are independent and identically scaled,
@@
yielding the familiar beta model. However, this construction does not explicitly exploit the intrinsic relationship between income distributions and inequality indices. In particular, the Gini coefficient can be expressed in terms of ratios of independent gamma variables, which suggests that suitable transformations of $X/(X+Y)$ may lead to probability models that are naturally tailored for inequality data.
In this work, we introduce two new families of unit–interval distributions defined by non–injective transformations of the ratio $X/(X+Y)$. Writing
$S_{r}(u)=[\,1-4u(1-u)\,]^{r},\qquad T_{r}(u)=1-[\,4u(1-u)\,]^{r},\quad 0<u<1,\ r>0,$
and setting $W=S_{r}\bigl(X/(X+Y)\bigr)$ and $Z=T_{r}\bigl(X/(X+Y)\bigr)$, one obtains random variables supported on $(0,1)$ whose densities admit simple closed–form expressions when $X$ and $Y$ are independent $\mathrm{Gamma}(\alpha_{i},\lambda_{i})$. The particular case $r=1/2$ recovers sample–based estimators of the Gini and Atkinson indices from a pair $(X,Y)$, thereby establishing a direct link between the proposed families and the classical measures of inequality (Gini,, 1936; Atkinson,, 1970). This connection provides a strong motivation for the use of the proposed models in the statistical analysis of Gini data.
The remainder of this paper is organized as follows. In Section 2, we introduce the new unit–interval families via the non–injective transformations $S_{r}$ and $T_{r}$ of the gamma ratio. In Section 3, we derive their distributional laws in terms of the underlying $X/(X+Y)$ density. In Section 4, we obtain closed–form expressions for the density and cumulative distribution functions of $W$ and $Z$. In Section 5, we present the quantile functions. In Section 6, we develop closed–form formulae for selected raw moments under special parameter regimes. In Section 7, we formulate maximum likelihood estimation, derive the likelihood equations, and discuss numerical solution strategies. In Section 8, we report the results of a Monte Carlo study. In Section 9, we illustrate the practical performance of the proposed models by fitting them to real Gini index data. Finally, in Section 10, we provide some concluding remarks. Additional technical derivations are collected in the Appendix.
@@ 2 The unit probabilistic models
respectively, where $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\exp(-t){\text{d}}t$, $z>0,$ is the (complete) gamma function. In this direction, a notable result in distribution theory, widely recognized in the literature, occurs when $\lambda_{1}=\lambda_{2}$ and $\varphi_{r}(x)=x,$ $0<x<1$, as this leads to ${X/(X+Y)}\sim\text{Beta}(\alpha_{1},\alpha_{2}).$ Alternative choices of $X$ and $Y$ beyond the gamma distributions have been thoroughly explored in the literature even for the case $\varphi_{r}(x)=x,$ $0<x<1$. For instance, Vila et al., 2024a selected $X$ and $Y$ as independent random variables with general distributions, whereas Vila et al., 2024c examined $X$ and $Y$ with a bivariate unit-logsymmetric distribution. Additionally, Vila et al., (2025) chose $X$ and $Y$ correlated according to the bivariate Birbaum-Saunders (BS) distribution, and Vila and Quintino, (2026) considered $X$ and $Y$ following a bivariate Fréchet distribution.
Simple examples include choosing the transformation $\varphi_{r}$ as injector functions, such as $\varphi_{r}(x)=x^{r},$ $\varphi_{r}(x)={2^{r}x^{r}/(x+1)^{r}},$ $\varphi_{r}(x)={(1-x)^{r}/(x+1)^{r}},$ $0<x<1$; $r>0$. The scope of this work centers on non-injective transformations $\varphi_{r}$, like:
@@
Note that both transformations above are symmetric and have a zero at the point $x=1/2$ and assume the value of $1$ at the points $x=0$ and $x=1$.
@@
$W\equiv S_{r}\left({X\over X+Y}\right)=\left({Y-X\over X+Y}\right)^{2r},\quad Z\equiv T_{r}\left({X\over X+Y}\right)=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}},$ (3)
@@
Remark $2.$
The value of $r=1/2$ in formulas (3) is special, because for this value the variables $W$ and $Z$ take the forms
$W={|X-Y|\over X+Y}\quad\text{and}\quad Z=1-{\sqrt{XY}\over{1\over 2}(X+Y)},$
@@ 3 Deriving the laws of $W$ and $Z$
$\mathbb{P}\left(W\in B\right)=\mathbb{P}\left(S_{r}\left({X\over X+Y}\right)\in B\right)$ $=\mathbb{P}\left(\left\{\omega\in\Omega:{X(\omega)\over X(\omega)+Y(\omega)}\in S_{r}^{-1}(B)\right\}\right)=\int_{S_{r}^{-1}(B)}f_{X\over X+Y}(u){\rm d}u,$ (4)
@@ 4 Distribution and density functions of $W$ and $Z$
Figures 1-3 display a range of shapes for the densities of $W$ and $Z$.
@@
It is worth noting that in the case $\lambda_{1}=\lambda_{2}\equiv\lambda$, the densities $f_{W}$ and $f_{Z}$ presented in Remark 4.1 do not depend on $\lambda$, consistent with the fact that $W$ and $Z$, given in (3), respectively, are derived from scale-invariant transformations. That is, for $X\sim\text{Gamma}(\alpha_{1},\lambda)$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda)$ we have $\lambda X\sim\text{Gamma}(\alpha_{1},1)$ and $\lambda Y\sim\text{Gamma}(\alpha_{2},1)$, and consequently, the random variables $W$ and $Z$ satisfy the identities
$W=\left({Y-X\over X+Y}\right)^{2r}=\left({\lambda Y-\lambda X\over\lambda X+\lambda Y}\right)^{2r},\quad Z=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}}=1-{(\lambda X\lambda Y)^{r}\over\displaystyle\left[{1\over 2}(\lambda X+\lambda Y)\right]^{2r}},$
@@ 6 Closed-form expressions for the $n$th raw moments of $W$ and $Z$
$\mathbb{E}(W^{n})={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}w^{n+{1\over 2r}-1}G(w_{r}){\rm d}w,\quad\mathbb{E}(Z^{n})={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}{z^{n}(1-z)^{{1\over r}-1}G(z_{r})\over\sqrt{1-(1-z)^{1\over r}}}{\rm d}z,$
where $w_{r}$ and $z_{r}$ are as given in (9) and (12), respectively, and $G(\cdot)$ is as in (16). Note that the above integrals are well-defined due to the support of $W$ and $Z$ in $(0,1)$, but numerical methods are necessary for their evaluation due to analytical complexity.
@@
6.1 The 11st raw moment of $W$: case $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$
Under the conditions $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$, in this part, we obtain (more precisely, in (23) and Remark 6.2) straightforward closed-form expressions for the 11st raw moment (expected value) of
@@ 6.1 The 11st raw moment of $W$: case $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$
$\mathbb{E}(W)=\mathbb{E}\left[\frac{|X-Y|}{X+Y}\right]=\mathbb{E}\left[|X-Y|\int_{0}^{\infty}\exp\left\{-\left(X+Y\right)x\right\}{\rm d}x\right]=\int_{0}^{\infty}\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]{\rm d}x,$ (18)
@@
$\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]$ $=\int_{0}^{\infty}\left[\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
@@
$\int_{0}^{\infty}\Bigg[\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\Bigg]\exp(-ux)=\mathbb{E}\left[\exp(-Xx)\right]\mathbb{E}\left[Y\exp(-Yx)\right]-\mathbb{E}\left[X\exp(-Xx)\right]\mathbb{E}\left[\exp(-Yx)\right]$
@@
$-\int_{0}^{\infty}\left[\int_{0}^{u}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$=2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$
@@
$+2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$ (20)
@@
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}.$ (21)
@@
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x.$ (22)
@@
$\mathbb{E}(W)$ $={\alpha_{2}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2}+1,\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)-{\alpha_{1}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2},\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}}{\rm d}x-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x$